### v1

# A Simple Knowledge Distillation Framework for Open-world Object Detection

Abstract
Open World Object Detection (OWOD) is a novel computer vision task with a considerable challenge, bridging the gap between classic object detection (OD) benchmarks and real-world object detection. In addition to detecting and classifying seen/known objects, OWOD algorithms are expected to localize all potential unseen/unknown objects and incrementally learn them. The large pre-trained vision-language grounding models (VLM, e.g., GLIP) have rich knowledge about the open world, but are limited by text prompts and cannot localize indescribable objects. However, there are many detection scenarios which pre-defined language descriptions are unavailable during inference. In this paper, we attempt to specialize the VLM model for OWOD task by distilling its open-world knowledge into a language-agnostic detector. Surprisingly, we observe that the combination of a simple knowledge distillation approach and the automatic pseudo-labeling mechanism in OWOD can achieve better performance for unknown object detection, even with a small amount of data. Unfortunately, knowledge distillation for unknown objects severely affects the learning of detectors with conventional structures for known objects, leading to catastrophic forgetting. To alleviate these problems, we propose the down-weight loss function for knowledge distillation from vision-language to single vision modality. Meanwhile, we decouple the learning of localization and recognition to reduce the impact of category interactions of known and unknown objects on the localization learning process. Comprehensive experiments performed on MS-COCO and PASCAL VOC demonstrate the effectiveness of our methods.
Introduction
Open-world object detection (OWOD) is a more practical detection problem in computer vision, facilitating the development of object detection (OD) (Ren et al. 2015; Zhu et al. 2020; Carion et al. 2020) in the real world. Within the OWOD paradigm, the model’s lifespan is pushed by an iterative learning process. At each episode, the model trained on the data with known object annotations is expected to detect known objects and all potential unknown objects. Human annotators then label a few of these tagged unknown classes of interest gradually. The model given these newly-added annotations continues incrementally updating its knowledge without retraining from scratch.
In the existing works (Joseph et al. 2021; Gupta et al. 2022; Yu et al. 2022; Wu et al. 2022a), the open-world detectors are expected to know about the open world through several datasets (Lin et al. 2014; Everingham et al. 2010) with tiny scales. However, the annotations of these datasets are too few to provide adequate object attributes for the model. It is difficult for the model to achieve the ideal goal through these datasets. The large pre-trained vision-language grounding models (VL) (Li et al. 2022; Zhang et al. 2022; Kamath et al. 2021; Radford et al. 2021; Yao et al. 2022) have rich knowledge of the open world due to countless millions of parameters, huge open-world datasets, and training costs.
However, their detection could not leave the participation of text prompts. Before detecting, the text prompt of all objects must be pre-listed to be detected, and the object whose text prompt was not listed could not be detected. There are many detection scenarios in which pre-defined language descriptions are unavailable during inference. Therefore, it poses a challenge how to equip the VLM with language-agnostic unknown object detection capability. In addition, their detection speed is also a criticism due to the following question. The huge number of parameters and FLOPs. The large pre-trained vision-language grounding model could only infer with several text prompts for the detecting performance, so they must infer many times when the number of prompts is large.
In this paper, we attempt to specialize the VLM model for OWOD task by distilling its open-world knowledge into a language-agnostic detector. We were surprised to observe that the combination of a simple knowledge distillation approach and the automatic pseudo-labeling mechanism in OWOD can achieve better performance for unknown object detection, even with a small amount of data. For the automatic pseudo-labeling mechanism, we set a box score prediction branch that leverages the predicted box score to automatically select pseudo-unknown labels from the remaining regression boxes after the matching process of each training iteration. So that the student could continuously evolve and learn the unknown beyond the large pre-trained vision-language grounding model. Unfortunately, knowledge distillation for unknown objects severely affects the learning of the detector with conventional structures for known objects due to the following challenges:
(I) Different from traditional Knowledge Distillation, the teacher and student contain different modalities and have different training manners, structures, and inference procedures. Therefore, the existing distilling objectives do not work and it is difficult to mix the learning of known grounding truth and unknown distilling knowledge. The performance of detecting known objects is also crucial for OWOD. Through the experiments, we reveal that the direct use of distilled knowledge dramatically affects the model’s learning ability of the original annotations. The model’s performance on detecting known objects is damaged substantially. To alleviate this, we propose the down-weight training loss for the distillation training, which utilizes the distilled labels’ object confidence from the large pre-trained vision-language grounding model to reduce the weight of unknown loss in the total loss during training.
(II) The presence of objects with highly similar features to known classes within the “unknown objects” can greatly affect the process of open-world object identification. This issue impacts not only the identification process but also the localization process for models that use coupled information for both tasks. Therefore, we propose decoupling the detection learning process. However, the decoupled structure completely separates the localization and identification, leading to many potential problems (e.g.mismatch between localization and identification results). Therefore, we propose a cascade structure that decouples the detecting process via two decoders and connects the two decoders via a cascade manner. In this structure, foreground localization can be protected from category knowledge because the loss of identification is diluted by the latter decoder. Moreover, the identification process can utilize the localization information because it leverages the former decoder’s output embeddings as the input queries.
Our contributions can be summarized fourfold:
We observe that the combination of the simple knowledge distillation and automatic pseudo-labeling mechanism could convert the open-world knowledge in the large pre-trained vision-language grounding model for the specialized OWOD task and propose a simple framework with surprisingly good performance.
To mitigate the effect of distilled knowledge on the detection performance of known objects, we propose the down-weight training loss function for the detector’s mixed learning process of known ground truth and distilled unknown knowledge.
We propose a cascade decoupled detection transformer structure to alleviate the influence caused by unknown objects on detecting known objects.
Our extensive experiments on two popular benchmarks demonstrate the effectiveness of our framework. Our model exceeds the distilled large pre-trained vision-language grounding model for OWOD and state-of-the-art methods for OWOD and IOD.
Related Works
Large pre-trained vision-language models: Recently, inspired by the success of vision-language(VL) pre-training methods (Radford et al. 2021) and their good zero-shot ability, (Gu et al. ; Li et al. 2022; Kamath et al. 2021; Zhang et al. 2022; Yao et al. 2022) attempted to perform zero-shot detection on a larger range of domains by using pre-trained vision language models. (Gu et al. ) proposed a zero-shot detection method to distill knowledge from a pre-trained vision language image classification model. (Li et al. 2022) tried to align region and language features using a dot-product operation and could be trained end-to-end on grounding and detection data. (Kamath et al. 2021) proposed an end-to-end modulated detector that detects objects in an image conditioned on a raw text query, like a caption or a question. (Yao et al. 2022) proposed a paralleled visual-concept pre-training method for open-world detection by resorting to knowledge enrichment from a designed concept dictionary.
Open-World Object Detection (OWOD): (Joseph et al. 2021) introduced OWOD task and ORE which adapted the faster-RCNN model with feature-space contrastive clustering, an RPN-based unknown detector, and an Energy Based Unknown Identifier (EBUI) for the OWOD objective. Recently, (Yu et al. 2022; Wu et al. 2022a, b) attempted to extend ORE. (Yu et al. 2022) was proposed to learn the discriminative embeddings of known classes in the feature space to minimize the overlapping distributions of known and unknown classes. (Wu et al. 2022a) proposed a two-branch objectness-centric open-world object detection framework consisting of the bias-guided detector and the objectness-centric calibrator. (Gupta et al. 2022) proposed to utilize the pseudo-labeling scheme to supervise unknown object detection, where unmatched object proposals with high backbone activation are selected as unknown objects.
Problem Formulation
denote the set of known object classes and denote the unknown classes which might be encountered at the test time, at the time . We labeled the known object categories in the dataset where denotes the input images and denotes the corresponding labels at time . The training image set consists of images and corresponding labels . Each denotes a set of object instances with their class labels and locations, denote the bounding box center coordinates, width and height respectively.
The artificial assumptions and restrictions in closed-set object detection are removed in Open-World Object Detection. It aligns object detection tasks more with real life. It requires the trained model to detect the previously encountered known classes and identify an unseen class instance as belonging to the unknown class. In addition, it requires the object detector to be capable of incremental updates for new knowledge, and this cycle continues over the detector’s lifespan. In the incremental updating phase, the unknown instances identified by are annotated manually. Along with their corresponding training examples, they update to and to . The model adds the new classes to known classes and updates itself to without retraining from scratch on the whole dataset .
Proposed method
This section elaborates on the proposed framework in detail. We start by illustrating the overall scheme of our proposed framework. Furthermore, we introduce the open-world object detector, the distillation process, and the matching and pseudo-labeling procedure. Last but not least, we describe the down-weight training strategy and inference phrase.
Overall Scheme
Fig.1 illustrates the overall scheme of our framework. For a given image , it is first sent into the open-world detector and the large pre-trained vision-language grounding model simultaneously. The detector leverages the visual features of the input to predict the localization, box score, and classification. The large pre-trained vision-language grounding model mines unknown open-world knowledge from the inputs. The known ground truth and unknown distilled knowledge aggregate the open-world supervision. In the training phase, we match the prediction and open-world supervision according to the regression loss, classification, and supervision confidence. After matching, the pseudo labels are selected according to the predicted box score. Then all labels are leveraged to train the open-world detector by the down-weight training loss function. In addition, when new categories are introduced at each episode, based on an exemplar replay-based finetuning to alleviate the catastrophic forgetting of learned classes and the finetuning by using a balanced set of exemplars stored for all known classes, our detector could continuously learn during its lifespan.
Open-World Object Detector
The open-world detector first uses a hierarchical feature extraction backbone to extract multi-scale features . The feature maps are projected from dimension to dimension by using 1×1 convolution and concatenated to vectors with dimensions after flattening out. Afterward, along with supplement positional encoding , the multi-scale features are sent into the deformable transformer encoder to encode semantic features. The encoded semantic features are acquired and sent into the localization decoder together with a set of learnable location queries. Aided by interleaved cross-attention and self-attention modules, the localization decoder transforms the location queries to a set of N location query embeddings . Meanwhile, the are used as class queries and sent into the identification decoder together with the (i.e.the encoded multi-scale features) again. The identification decoder transforms the class queries to class query embeddings that correspond to the location query embeddings. The operation of cascade decoders is expressed as follows:
where and denote the localization and identification decoder. is the encoder and is the backbone. represents the reference points and denotes the input image. stands for the class queries. Eventually, the are then sent into the classification branch to predict the category . The are then input to the regression and box score branch to locate N foreground bounding boxes and predict the box score .
Open-World Object Supervision
In the knowledge distillation phase, the teacher leverages a large pre-trained vision-language grounding model and a text prompt that contains object categories as many as possible to mine unknown open-world knowledge from input . In this paper, we utilize GLIP (Li et al. 2022) and categories of LVIS (Gupta, Dollar, and Girshick 2019) as the large pre-trained vision-language grounding model and the text prompts, respectively. The distilled open-world knowledge is first processed through the NMS produce. Then we align the distilled open-world knowledge and ground truth by the align module, where we align the generated labels to the given data annotation space e.g.translate LVIS categories (trailer truck, tow truck, et al.) into COCO category truck and exclude the known set in the distilled open-world knowledge. In addition, the identification confidence of the unknown labels from the large pre-trained vision-language grounding model is reserved for the following training process and leveraged as the supervision confidence.
Matching and Evolving
When matching the open-world supervision and prediction, we consider the box regression loss, prediction class score, and confidence of supervision. After matching, we leverage the box score branch to help the detector learn more unknown open-world objects beyond the ground truth and distilled knowledge from the large pre-trained vision-language grounding model during training via selecting pseudo labels. We denote by the open-world supervision, and the set of predictions. For finding the best bipartite matching between them, a permutation of elements with the lowest cost is searched for as follows:
where is the pair-wise matching cost between which represents ground truth or labels from the large pre-trained vision-language grounding model and a prediction with index , shown as Equation.4. Inspired by (Zhu et al. 2020; Carion et al. 2020; Kamath et al. 2021), we choose the Hungarian algorithm as the optimal assignment.
where denotes the regression loss, which consists of box loss and GIOU loss (Rezatofighi et al. 2019). and represent the open-world supervision box and prediction box, respectively. is the confidence of the open-world supervision. Then, the pseudo labels are selected as follows:
where denotes the prediction box score. The pseudo-labels could prevent the model from falling entirely into the knowledge of the large pre-trained vision-language grounding model and help it know unseen objects beyond the distilled knowledge.
| Task IDs | Task 1 | Task 2 | Task 3 | |||
| Metrics | Unknown | Known | Unknown | Known | Unknown | Known |
| Recall () | mAP() | Recall() | mAP() | Recall() | mAP() | |
| Baseline(GLIP) | 37.0 | 42.0 | 32.4 | 32.4 | 34.9 | 28.2 |
| Distillation | 39.4 | 41.1 | 36.6 | 19.2 | 36.6 | 14.1 |
| Distillation + DW | 39.2 | 53.5 | 36.3 | 34.9 | 35.9 | 25.1 |
| Distillation + CS | 39.1 | 51.9 | 36.9 | 34.6 | 36.1 | 25.9 |
| Final: Ours | 39.0 | 56.8 | 36.7 | 40.3 | 36.1 | 30.1 |
| Prompt | SPLIT | Method | mAP | UR | UR(sam) |
| GLIP | 42.2 | 37.0 | 8.04 | ||
| OWOD | Ours | 56.8 | 39.0 | 16.5 | |
| GLIP | 46.5 | 52.6 | 9.4 | ||
| LVIS | MSCOCO | Ours | 69.4 | 60.9 | 20.0 |
| GLIP | - | - | 7.5 | ||
| OWOD | Ours | 57.6 | 38.8 | 16.8 | |
| GLIP | - | - | 7.6 | ||
| LVIS w/o COCO | MSCOCO | Ours | 70.6 | 58.9 | 20.4 |
| GLIP | 67.8 | 43.3 | 3.2 | ||
| OWOD | Ours | 60.9 | 41.1 | 12.9 | |
| GLIP | 73.1 | 64.8 | 5.3 | ||
| COCO | MSCOCO | Ours | 74.8 | 69.6 | 16.7 |
| SPLIT | Metrics | DDETR | CAT | Ours |
| mAP | 53.5 | 55.3 | 56.8 | |
| UR | 39.2 | 39.4 | 39.0 | |
| OWOD | UR(sam) | 16.3 | 16.5 | 16.5 |
| Metric | ||||
| Teacher | SPLIT | mAP | UR | UR(sam) |
| GLIP | OWOD | 56.8 | 39.0 | 16.5 |
| SAM | OWOD | 19.4 | 33.5 | 38.3 |
Down-Weight Training Strategy
The inclusion of distilled knowledge inevitably influences the model’s learning of the known set. Because the quality of it could not be guaranteed, and it increases the difficulty of detector learning. Therefore, we propose the down-weight training strategy, which leverages the distilled knowledge identification confidence to generate harmonic factor and down-weight the unknown training loss and train the detector in an end-to-end manner as shown in Fig.1 (b). The training loss function is as follows:
where the uses the common regression loss which consists of box and GIOU loss (Rezatofighi et al. 2019). and represent the box score and classification loss, respectively. They all leverage the common sigmoid focal loss (Lin et al. 2017). For simplicity, we omit them. In addition, the , , and all utilize correspondingly down-weight loss function we propose, the formulations are shown as follows:
where and denote the distilled knowledge supervision and pseudo supervision labels, respectively. represents the number of queries, represents the index of the label corresponding to the prediction. denotes the sigmoid focal loss function. represents the GIOU loss function. is the confidence of supervision labels. , , and are the prediction box, box score, and classification score, respectively.
Inference
During inference, our detector only utilizes the visual features of the inputs to detect open-world objects without any information from the other modalities. The inference process is to composite the detector output to form open-world object instances. Formally, the -th output prediction is generated as . According to the formulation : , the result is acquired as , where is the category, denotes the confidence, and represents the predicted bounding boxes.
Experiment
Datasets and Metrics
For a fair comparison, we implement the experiments on two mainstream splits of MS-COCO (Lin et al. 2014), and Pascal VOC (Everingham et al. 2010) dataset. We group the classes into a set of non-overlapping tasks . The class in task only appears in tasks where . In task , classes encountered in and are considered as known and unknown classes, respectively.
OWOD SPLIT (Joseph et al. 2021) splits the 80 classes of MS-COCO into 4 tasks and selects a training set for each task from the MS-COCO and Pascal VOC training set. Pascal VOC testing and MS-COCO validation set are used for evaluation.
MS-COCO SPLIT (Gupta et al. 2022) mitigates data leakage across tasks in (Joseph et al. 2021) and is more challenging. The training and testing data are selected from MS-COCO.
Metrics: Following the most commonly used evaluation metric for object detection, we use mean average precision (mAP) to evaluate the known objects. Following OW-DETR (Gupta et al. 2022), U-Recall which measures the ability of the model to retrieve unknown object instances for OWOD problems, is used as the metric for unknown objects.
Implementation Details
The multi-scale feature extractor consists of a Resnet-50(He et al. 2016) pretrained on ImageNet(Deng et al. 2009) in a self-supervised(Caron et al. 2021) manner and a deformable transformer encoder whose number of layers is set to 6. For the two cascade decoders, we all use the deformable transformer decoder, and the number of layers is set to 6, too. We set the number of queries , the dimension of the embeddings , and the number of pseudo-labels . We set GLIP (Li et al. 2022) as the large pre-trained vision-language grounding model and categories of LVIS dataset (Gupta, Dollar, and Girshick 2019) as text prompts to assist the training process. For GLIP, we use the GLIP-L (Li et al. 2022) without finetuning on the COCO dataset (Lin et al. 2014). It consists of Swin-Large (Liu et al. 2021), text encoder of CLIP (Radford et al. 2021), DyHead (Dai et al. 2021), BERT Layer (Devlin et al. 2018) and Fusion Module (Li et al. 2022). For the incremental object detection experiments, we only use our open-world detector without the help of GLIP.
Ablation Study
In this subsection, a set of experiments are designed to clearly understand the contribution of each of the constituent components. We conducted all experiments on the OWOD split.
| Task IDs | Task 1 | Task 2 | Task 3 | |||
| Inference | Unknown | |||||
| Param# | Rate | Recall() | ||||
| SPLIT | Metrics | (s/img) | ||||
| OWOD | GLIP | 321.9M | 9.22 | 37.0 | 35.5 | 34.9 |
| Ours | 42.9M | 0.08 | 39.0 | 36.7 | 36.1 | |
| MS-COCO | GLIP | 321.9M | 9.22 | 52.6 | 54.5 | 53.3 |
| Ours | 42.9M | 0.08 | 60.9 | 60.0 | 58.6 | |
Ablating components: To study the contribution of each component, we design ablation experiments in Table.1 (a). We set GLIP as our baseline, the Distillation improves the performance on unknown object detection but reduce the detector’s ability for known objects. Adding the down-weight training loss function significantly improves the performance on detecting known objects and incremental object detecting, achieving significant absolute gains up to more than 10 points. As we have analyzed, the cascade decoupling structure alleviates the inclusion of unknown objects on the known detecting performance and the confusion between the categories and locations of the same objects. It significantly improves the performance of detecting known objects with absolute gains up to more than 10 points, too. What’s more, these two combine effectively, improving performance with absolute gains up to 15.7, 21.1 and 16.0, without reducing the ability to detect unknown objects. Thus, each component has a critical role to play in open-World object detection.
Ablating different prompts: To ablate the impact of text prompts for knowledge distillation, we do experiments on different text prompts, i.e.LVIS, LVIS without COCO and COCO as shown in Table.1(a). From the experiment results (i.e.UR comparison between LVIS and COCO prompt and qualitative results in Table.1(d)), we investigate that the unknown annotations in the test set are not adequate and more corresponding to the original 80 categories in COCO dataset. Therefore, existing “unknown recall” can not evaluate the detecting ability for unknown open-world objects beyond COCO annotation. To evaluate that, we use SAM(SAM) to generate annotations for all areas of similar objects in the test set. Though these annotations are noisy and contain many non-object box annotations, the recall for these can evaluate the unknown detection ability of detectors. The comparison between LVIS and COCO prompts shows that less distilled knowledge reduces the impact on learning known objects. In addition, less distilled knowledge impacts the performance of the knowledge distillation framework. As shown in OWOD split of COCO prompt, our detector does not learn better detecting performance for the original unknown objects. The results in LVIS w/o COCO prompt demonstrate our framework letter more unknown open-world objects. Through we exclude all corresponding coco categories from the LVIS prompt, our detector’s detection ability for original unknown objects does not seem to be impacted.
Ablating different detection transformer structures: As shown in Table.1(b), we compare our structure with original parallel structure (Zhu et al. 2020) and structure of (Ma et al. 2023). The results demonstrate that our structure has a better ability for alleviating the influence of distilled knowledge on known objects.
Ablating different teachers: In Table.1(c), we compare different teachers, i.e.GLIP(Li et al. 2022) and SAM (SAM). For SAM, due to the absence of supervision confidence, we can not leverage the down-weight loss function. Therefore, the known detection ability is severely reduced. In addition, due to the over-detection ability of SAM, the detector can not distill a better unknown detection ability.
Ablating the Pseudo Labels: We also ablate the pseudo labels by controlling the presence or absence of the box score branch. However, owing to the absence of effective metrics, existing metrics cannot demonstrate the difference. In terms of Unknown Recall, they almost have the same performance on UR and UR(sam) when the box score branch is removed (i.e.While it is possible to add a lot of unknown annotations after adding SAM-generated annotations, there are too many non-object annotations. Also, they can not evaluate the performance of unknown object detection in absolute terms). Therefore, we ablate it according to the several visualization results, as shown in Fig.2. When the pseudo labels are removed, the model misdetect many unknown objects. The comparison demonstrates that the pseudo labels can help the detector better know about the world.
| Task IDs | Task 1 | Task 2 | Task 3 | Task 4 | ||||||||||
| Unknown | Known | Unknown | Known | Unknown | Known | Known | ||||||||
| Recall | mAP() | Recall | mAP() | Recall | mAP() | mAP() | ||||||||
| Metrics | Split | () | Current | () | Previously | Current | Both | () | Previously | Current | Both | Previously | Current | Both |
| UC-OWOD | OWOD | 2.4 | 50.7 | 3.4 | 33.1 | 30.5 | 31.8 | 8.7 | 28.8 | 16.3 | 24.6 | 25.6 | 15.9 | 23.2 |
| ORE-EBUI | 4.9 | 56.0 | 2.9 | 52.7 | 26.0 | 39.4 | 3.9 | 38.2 | 12.7 | 29.7 | 29.6 | 12.4 | 25.3 | |
| OW-DETR | 7.5 | 59.2 | 6.2 | 53.6 | 33.5 | 42.9 | 5.7 | 38.3 | 15.8 | 30.8 | 31.4 | 17.1 | 27.8 | |
| OCPL | 8.3 | 56.6 | 7.7 | 50.6 | 27.5 | 39.1 | 11.9 | 38.7 | 14.7 | 30.7 | 30.7 | 14.4 | 26.7 | |
| 2B-OCD | 12.1 | 56.4 | 9.4 | 51.6 | 25.3 | 38.5 | 11.6 | 37.2 | 13.2 | 29.2 | 30.0 | 13.3 | 25.8 | |
| Ours | 39.0 | 56.8 | 36.7 | 52.3 | 28.3 | 40.3 | 36.1 | 36.9 | 16.4 | 30.1 | 31.0 | 14.7 | 26.9 | |
| ORE-EBUI | MS-COCO | 1.5 | 61.4 | 3.9 | 56.5 | 26.1 | 40.6 | 3.6 | 38.7 | 23.7 | 33.7 | 33.6 | 26.3 | 31.8 |
| OW-DETR | 5.7 | 71.5 | 6.2 | 62.8 | 27.5 | 43.8 | 6.9 | 45.2 | 24.9 | 38.5 | 38.2 | 28.1 | 33.1 | |
| Ours | 60.9 | 69.4 | 60.0 | 63.8 | 26.9 | 44.4 | 58.6 | 46.2 | 28.0 | 40.1 | 41.8 | 29.6 | 38.7 | |
| Method | Settings | ||
| 10+10 | 15+5 | 19+1 | |
| ILOD | 63.2 | 65.8 | 68.2 |
| Faster ILOD | 62.1 | 67.9 | 68.5 |
| ORE | 64.5 | 68.5 | 68.8 |
| OW-DETR | 65.7 | 69.4 | 70.2 |
| Ours | 68.6 | 72.1 | 72.9 |
More Detailed Comparison With the Teacher
In order to keep the model’s ability to generalize to open-world objects in the comparison, we do not finetune the teacher on the known dataset. Because the finetuning process severely destroy the teacher’s generality. Although we also show the detection accuracy for the known classes, we only need to focus on the comparison of the detection ability for unknown objects. The results compared with the large pre-trained vision-language grounding model (GLIP) for OWOD problem are shown in Table.2. Regarding the number of parameters and FLOPs, our model is significantly smaller than GLIP. In particular, the inference speed of ours is of GLIP’s. Furthermore, the number of GLIP’s training data is 64M images, while our model only needs a small amount of data in each task of different splits almost of its. Compared with GLIP’s U-Recall of 37.0, 35.5, and 34.9 on Task 1, 2, and 3 of OWOD split, ours achieves 39.0, 36.7 and 36.1 in the corresponding tasks, achieving significant absolute gains up to 2.0, 1.2 and 1.2, respectively. In the MS-COCO split, our model achieves more than 5 point improvement in terms of U-Recall. This demonstrates that our model has a better ability to detect unknown objects for OWOD.
Qualitative Results: To present the ability to detect unknown objects intuitively, we select several open scenes from games and comics. These scenes contain the object out of the categories of LVIS (Gupta, Dollar, and Girshick 2019) text prompts. visualization results are shown in Fig.3. Benefits from the box score branch and pseudo labels, our detector evolves the novel unknown objects beyond the large pre-trained vision-language grounding model.
Comparison With SOTA OWOD Methods
OWOD SPLIT: The results compared with the state-of-the-art methods on OWOD split for OWOD problem are shown in Table.3 (a). The ability of our model to detect unknown objects quantified by U-Recall is more than of those reported in previous state-of-the-art OWOD methods. Compared with the model 2B-OCD (Wu et al. 2022a) with the highest U-Recall of 12.1, 9.4 and 11.6 on Task 1, 2, and 3, ours achieves 39.0, 36.7, and 36.1 in the corresponding tasks, achieving significant absolute gains up to 26.9, 27.3 and 24.5, respectively. Benefits from the cascade decoder structure and down-weight training loss which mitigate the effect of unknown objects on detecting known objects, our model’s performance on known objects is also superior to most state-of-the-art methods.
MS-COCO SPLIT: We report the results on the MS-COCO split in Table.3 (a). MS-COCO split mitigates data leakage across tasks and assigns more data to each Task, while our model receives a more significant boost compared with the OWOD split. Our model’s unknown object detection capability, quantified by U-Recall, is almost of those reported in previous state-of-the-art OWOD methods. Compared with OW-DETR’s U-Recall of 5.7, 6.2, and 6.9 on Task 1, 2, and 3, ours achieves 60.9, 60.0, and 58.6 in the corresponding tasks, achieving significant absolute gains up to 55.2, 53.8 and 51.7, respectively. This demonstrates that our model has the more powerful ability to retrieve new knowledge and detect unknown objects when faced with more difficult tasks.
Incremental Object Detection: To intuitively present our detector’s ability for detecting object instances, we compare it to (Shmelkov, Schmid, and Alahari 2017; Peng, Zhao, and Lovell 2020; Joseph et al. 2021; Gupta et al. 2022) on the incremental object detection (IOD) task. We do not use assistance from the large pre-trained vision-language grounding model. We evaluate the experiments on three standard settings, where a group of classes (10, 5, and last class) is introduced incrementally to a detector trained on the remaining classes (10, 15, and 19), based on the PASCAL VOC 2007 dataset (Everingham et al. 2010). As the results shown in Table.3 (b), our model outperforms the existing method in a great migration on all three settings, indicating the power of cascade detection transformer for IOD.
Conclusions
In this paper, we attempt to specialize the VLM model for OWOD task by distilling its open-world knowledge into a language agnostic detect. we observe that the combination of the simple knowledge distillation and automatic pseudo-labeling mechanism could convert open-world knowledge. To mitigate the effect of distilled knowledge on the detection performance of known objects, we propose the down-weight training loss function for the detector’s mixed learning process of known ground truth and distilled unknown knowledge. We propose a cascade decoupled detection transformer structure to alleviate the influence caused by unknown objects on detecting known objects. Our extensive experiments on two popular benchmarks demonstrate the effectiveness of our framework.
Acknowledgment
This work is supported by National Nature Science Foundation of China (grant No.61871106), and the Open Project Program Foundation of the Key Laboratory of Opto-Electronics Information Processing, Chinese Academy of Sciences (OEIP-O-202002).